OpenStudy (anonymous):

Explain, in complete sentences, how a higher order root is different from a square root. Give two examples.

hero (hero):

3*3 = 9 sqrt{9} = 3 3*3*3 = 9*3 = 27 cubert{27} = 3 If a number is squared, you can take the square root of a square to get back to the original number. Square root is the inverse of square. If a number is cubed, you can take the cube root of a cube to get back to the original number. Cube root is the inverse of a cube.

OpenStudy (anonymous):

thank you (:

hero (hero):

yw